OpenStudy (anonymous):

For which operations is the set {–1, 1} closed? Choose all answers that are correct. A. addition B. subtraction C. multiplication D. division

OpenStudy (zzr0ck3r):

-1+1 = 0 is not in the ser -1-1 = -2 is not in the ser 1*1=-1*-1=1 is in the set -1*1=1*-1=-1 is in the set similarly for division